Find the following limits


lim( sqrt2n+1− Sqrt2n)

n→∞


lim3n^3 −n+8/(4n(n−1)(n−2))

n→∞



b) Consider the sequence (an) = (−1)^n − 2n

 Explain whether the sequence is monotone increasing or decreasing, whether it is monotone and if limn→∞(an) = −∞



1
Expert's answer
2021-05-13T08:37:30-0400

(a) (i)"lim_{n\to \infty} (\sqrt{2n+1}-\sqrt{2n})"


          Rationalising the given expression-

                

        "= lim_{n\to \infty} (\sqrt{2n+1}-\sqrt{2n})\times \dfrac{ \sqrt{2n+1}+\sqrt{2n}}{\sqrt{2n+1}+\sqrt{2n}} 



 \\[9pt]

 = lim_{n\to \infty} \dfrac{ (\sqrt{2n+1})^2-(\sqrt{2n})^2}{\sqrt{2n+1}+\sqrt{2n}} 


\\[9pt]
 =lim _{n\to \infty} \dfrac{1}{\sqrt{2n+1}+\sqrt{2n}}"

        Applying limits


         = 0


(ii) "lim_{n\to \infty} \dfrac{3n^3-n+8}{4n(n-1)(n-2)}"

 

    Now writing it my simplifying-

     

   "=lim_{n\to \infty} \dfrac{3n^3-n+8}{4n(n^2-2n-n+2)} 

 \\[9pt]

 =lim_{n\to \infty} \dfrac{3n^3-n+8}{4n^3-12n^2+8n}



 \\[9pt]=lim_{n\to \infty} \dfrac{n^3(3-\frac{1}{n^2}+\frac{8}{n^3})}{n^3(4-\frac{12}{n}+\frac{8}{n^2})}



 \\[9pt]=\dfrac{3-0+0}{4-0+0}=\dfrac{3}{4}"

    


(b) Given, sequence

        "a_n=(-1)^n-2n"

        "a_1=-1-2=-3\\

 a_2=1-4=-3\\

 a_3=-1-6=-7\\

 a_4=-7\\

 a_5=-1-10=-11\\

 a_6=1-12=-11\\"


"\Rightarrow <a_n>=-3,-3,-7,-7,-11,-11..."


 Here, "a_1\ge a_2\ge a_3\ge a_4\ge a_5\ge a_6..."


we know a sequence "<a_n>" is monotonic decreasing if "a_{n+1}\le a_n \forall n\in N"


Hence Given sequence is monotone decreasing.

So "lim_{n\to \infty}(a_n)=-\infty."